# Quantum Rainbow Codes

### v2

@@ II.2 The simplex graph
From Proposition 1 we can see that an $X$ operator supported on the vertices of an $x$-rainbow subgraph commutes with all $Z$ stabilisers (and similarly for $Z$ operators on $z$-rainbow subgraphs) and so if these operators are not stabilisers they must be logicals. We will see in Section III that this fact is the cause of the low distances observed for the majority of pin codes studied numerically in vuillot2022quantum , as the method used to generate these codes naturally gives rise to rainbow subgraphs of size $4$. In addition, in Section IV we will show that it prevents useful transversal non-Clifford gates in the majority of cases.
@@ II.3 Rainbow codes
A $D$-dimensional rainbow code is a CSS code with qubits associated to the vertices of a simplex graph with $D+1$ colours and stabiliser generators associated to a subset of the $x$-maximal and $x$-rainbow subgraphs (for $X$ stabilisers) and $z$-maximal and $z$-rainbow subgraphs (for $Z$ stabilisers) of this graph, for $x$ and $z$ satisfying Eq. 1.
@@
Note that in general mixed rainbow codes are not required to have $X$ and $Z$ stabilisers supported on all $x$- and $z$-rainbow subgraphs as such operators would not commute in general. The definitions of these different classes of rainbow code are summarised in Table 1. Notice that all of these classes can be thought of as generalisations of colour codes as maximal and rainbow subgraphs in colour codes are equivalent. To understand how the properties of these classes can differ from each other we will need to look at some more specific examples, such as those presented in the next section. We also remark that, for all stabiliser assignments we will consider, the sets of all relevant maximal and rainbow subgraphs are efficiently computable and algorithms for finding them are presented in Appendix C.
@@
mixed some $x$-maximal and $x$-rainbow some $z$-maximal and $z$-rainbow
@@ III Rainbow codes from hypergraph products and gluings of colour codes
Given a chain complex that defines a cellulation of a manifold, a colour code on this manifold can be defined using a mapping from the chain complex to a corresponding simplicial complex bombin2007exact ; zhu2023non . Similar techniques were used in vuillot2022quantum to define pin codes on more general chain complexes obtained from the hypergraph product of classical codes tillich_quantum_2014 , and we could use these mappings and the results of the previous section to obtain other classes of rainbow code in an identical fashion. However, it turns out that there is an alternative perspective on the hypergraph product – and on the codes derived from it – that provides considerably more intuition into the structure and properties of these codes. In what follows we will show that any $D$-dimensional rainbow code defined via the hypergraph product can be viewed as a collection of $D$-dimensional colour codes joined together at $(D-1)$-dimensional domain walls. This makes understanding the code parameters and action of logical gates much more straightforward, and also implies connections between these codes and other constructions involving some forms of code gluing, such as welded codes michnicki20143d , defect networks song2023topological , and layer codes williamson_layer_2024 .
We begin by studying gluing operations on graphs and their interplay with the hypergraph product in Section III.1. In Section III.2 we show how to obtain flag graphs from the output of the hypergraph product and demonstrate the effects of gluing operations on these graphs. In Section III.3 we show how, when all inputs to the product are cycles, the resulting flag graph can be understood as a collection of colour code lattices joined at objects we call seams, and finally we show how to assign stabilisers to these seams and study their effect on colour code logical operators in Section III.4.
@@ III.1 Gluings of graphs
To start, we define the hypergraph product (HGP) and show how its action commutes with gluing operations on graphs. This product is commonly defined in terms of tensor products of chain complexes but it can also be understood as a Cartesian product of bipartite graphs  [^2] If we label the two kinds of vertex type-00 and type-$1$ then the adjacency matrix with type-00 vertices as rows and type-$1$ vertices as columns is equivalent to a representation of the boundary map from $1$-cells to 00-cells in the chain complex picture. Recall that a graph $\mathcal{G}$ is a set of vertices $V_{\mathcal{G}}$ and a set of edges $E_{\mathcal{G}}$, where an edge is an (unordered) pair of vertices $\{g_{1},g_{2}\}$. Also, the Cartesian product of two graphs is defined as
@@
Given a graph $\mathcal{G}\square\mathcal{H}$, the operation $g_{1}\Leftarrow g_{2}:\mathcal{G}\square\mathcal{H}$ is a gluing $((g_{1},h_{i})\leftarrow(g_{2},h_{i}):\mathcal{G}\square\mathcal{H})\leavevmode\nobreak\ \forall\leavevmode\nobreak\ h_{i}\in V_{\mathcal{H}}$
where the order of these gluings does not matter by Lemma 4. This operation generalises straightforwardly to gluing along a hyperplane in a graph $\mathcal{G}=\mathcal{A}\square\mathcal{B}\square...$ since we can use the associativity and commutativity of the Cartesian product to write
$\mathcal{G}=\mathcal{F}\square(\mathcal{A}\square\mathcal{B}\square...)=\mathcal{F}\square\mathcal{H}$ (5)
@@
$\{\{g_{i},h_{j}\}\leavevmode\nobreak\ |\leavevmode\nobreak\ (g_{i}\neq g_{2})\in V_{\mathcal{G}},h_{j}\in V_{\mathcal{H}}\}.$ (8)
On the other hand, $g_{1}\Leftarrow g_{2}:\mathcal{G}$ results in a vertex set $V_{\mathcal{G}}^{\prime}=\{g_{i}\leavevmode\nobreak\ |\leavevmode\nobreak\ (g_{i}\neq g_{2})\in V_{\mathcal{G}}\}$ and the product $(g_{1}\Leftarrow g_{2}:\mathcal{G})\square\mathcal{H}$ then has vertex set
$\{\{g_{i},h_{j}\}\leavevmode\nobreak\ |\leavevmode\nobreak\ g_{i}\in V_{\mathcal{G}}^{\prime},h_{j}\in V_{\mathcal{H}}\}$ (9)
@@
Secondly, we can show that each vertex has the same neighbourhood in both graphs (equivalent to showing that the edge sets are equivalent). First, notice that for a vertex $(g_{i},h_{j})$ of $\mathcal{G}\square\mathcal{H}$, the neighbourhood $\mathcal{N}((g_{i},h_{j}))$ can be split into two parts, $N^{g}$ and $N^{h}$, where $N^{g}=\mathcal{N}(g_{i})\times h_{j}$ and $N^{h}=g_{i}\times\mathcal{N}(h_{j})$. Then, recalling the effect of gluing on neighbourhoods described in the proof of Lemma 4, for $g_{1}\Leftarrow g_{2}:\mathcal{G}\square\mathcal{H}$ we have that
@@
An example of this commutation is shown in Fig. 3.
@@
[…]
@@
[…]
@@
[…]
is then guaranteed by Proposition 2 and Eq. 14.
The insight provided by Eq. 16 that instead of studying the product of a few very complex graphs, we can equivalently “unglue” these graphs into simpler graphs, take products of these and then glue together the results. However, we still need a way to define simplicial complexes from these graphs, as well as way to understand the effect of gluing on these complexes.
@@ III.2 Flags and their gluings
More generally, for the Cartesian product of $D$ bipartite graphs we will have $D+1$ levels of vertex in the output, with the level of a given output vertex equal to the sum of the levels of all its corresponding input vertices. Typically we will be focused on a single input graph $\mathcal{G}$ (i.e. as the target of a gluing) and in this case we will use an abbreviated notation $(g_{i}^{l},...)^{L}$ where $l\in\{0,1\}$, $L$ is the level of the product vertex and the ... represents input vertices from all other input graphs. A $D$-dimensional simplicial complex can then be obtained from this graph by associating a $D$-simplex with every path of length $D+1$ that consists of a sequence of vertices with levels $0,1,...,D$. In keeping with the terminology of vuillot2022quantum we will refer to these paths as flags and write them as $(D+1)$-uples $((g_{i}^{0},...)^{0},...,(g_{i^{\prime}}^{1},...)^{D})$. Two examples for $D=2$ and $D=3$ are shown in Fig. 4. When discussing simplicial complexes obtained in this manner we will use the term “flag graph” (or $\mathcal{G}_{\mathcal{F}}$) to refer to the simplex graph and “product graph” (or $\mathcal{G}_{\square}$) to refer to the graph with levelled vertices obtained from the Cartesian product. Additionally, the vertex levels in the product graph will be used to label the edge colours in the flag graph, so that e.g. two flags which share product graph vertices with levels $1,2,...,D$ will correspond to flag graph vertices connected by an edge with colour $c_{0}$.
@@
Figure 5: Gluings on product graphs (numbered vertices and black edges) and their effects on the corresponding flag graphs (dark vertices and coloured edges). a) A single gluing that joins two flags at $D-1$ vertices in the product graph and so adds a new $c_{0}$-coloured edge between the corresponding vertices of the flag graph. b) A pair of gluings that glue together pairs of flags in the product graph and so glue together the corresponding vertices in the flag graph. Seams in the product graph are shown in both cases by grey edges/vertices.
When considering gluings of bipartite graphs, bipartiteness is only preserved when gluings are performed between vertices of the same level, so we will consider only these types of gluings. The flag graph is modified by gluings of the product graph only when flags in the product graph are glued together at $D$ or $D+1$ of their vertices, with the former creating new edges in the flag graph and the latter gluing together vertices of the flag graph. Notice, however, that this second case cannnot be accomplished with a single hyperplane gluing as, for example, the gluing $g_{i}^{0}\Leftarrow g_{j}^{0}$ can non-trivially affect at most $D$ vertices of any flag, as a flag where all $D$ vertices are of the form $(g_{i}^{0},...)$ or $(g_{j}^{0},...)$ would not contain a level $D$ vertex. On the other hand, we can identify two distinct cases where flags can be joined at $D$ vertices by a single gluing. Either we have a gluing $g_{i}^{0}\Leftarrow g_{j}^{0}$ and a pair of flags
@@
In the first case a new $c_{D}$ edge in the flag graph will be created between the vertices corresponding to $F_{1}$ and $F_{2}$, while in the second case a new $c_{0}$ edge will be created. No other colours of edge can be created by gluing. An example of such a gluing is shown in Fig. 5 a). We then fix the following terminology.
@@
As alluded to above, it is possible to use multiple gluings to glue one flag fully onto another, and thus glue together a pair of vertices in the flag graph. This is accomplished using a pair of gluings $g_{i}^{0}\Leftarrow g_{j}^{0}$ and $g_{i^{\prime}}^{1}\Leftarrow g_{j^{\prime}}^{1}$ where $g_{i^{\prime}}^{1}\in\mathcal{N}(g_{i}^{0})$ and $g_{j^{\prime}}^{1}\in\mathcal{N}(g_{j}^{0})$. This creates a pair of seams, as in Fig. 5 b).
[…]
@@ III.3 Joining colour code lattices
Because $\mathcal{G}_{\square}$ is a cellulation of a $D$-dimensional manifold the simplicial complex described by $\mathcal{G}_{\mathcal{F}}$ is also a cellulation of this same manifold (it is simply a subdivision of the hypercubic lattice into simplices), and by Lemma 1 $\mathcal{G}_{\mathcal{F}}$ is then a $D$-dimensional colour code lattice. ∎
@@
An example of this is shown in Fig. 6 a). This statement is equivalent to results in other works where $\mathcal{G}_{\square}$ is interpreted as a chain complex rather than a graph bombin2007exact ; vuillot2022quantum .
@@
Now consider a graph $\mathcal{G}_{\square}=(\mathcal{O}_{1}\cup\mathcal{O}_{2})\square\mathcal{O}_{3}\square...\square\mathcal{O}_{D+1}$ which is equivalent to a pair of disjoint $D$-dimensional hypercubic lattices, and so has a flag graph equivalant to a pair of disjoint $D$-dimensional colour code lattices. The gluing $o_{1,i}^{l}\Leftarrow o_{2,j}^{l}:\mathcal{G}_{\square}$ which glues together these two hypercubic lattices then also modifies the associated colour code lattices/flag graph in accordance with the discussion in the previous section. Specifically, it creates new $c_{0}$ or $c_{D}$ edges (depending on $l=1$ or 00) between vertices of $\mathcal{G}_{\mathcal{F}}$ that correspond to flags lying on the seam assocaited to the gluing, and so joins the whole flag graph into a single connected component. An example is shown in Fig. 6 b), where we can see that the flag graph post-gluing remains locally identical to a colour code lattice everywhere except in the neighbourhood of the seam. We refer to this operation as the join of two colour code lattices.
@@
[…]
Notice that, due to Eq. 15, it is not guaranteed that performing a join and then a split recovers the original pair of colour code lattices. However, the requirement that $\lvert U\rvert=2$ ensures that we recover either a pair of colour code lattice or a single larger colour code lattice, as this ungluing necessarily divides the underlying glued cycle graphs back into a disjoint union of cycle graphs. We can also convince ourselves of this diagramatically by considering a gluing of two cycles
@@
Sets of colour code lattices that can be mapped into each other by performing a join and then a split will be referred to as split-equivalent. Examples are shown in Fig. 6 c) and d).
@@
where two cycles are glued at three vertices/two edges. Because two of the vertices in the resulting graph have degree three there is no way to recover a pair of cycles using the ungluing operation that we have defined. We show in Fig. 7 the effect of such a gluing on a pair of colour code lattices. Because the gluing is not reversible there is no splitting of this flag graph that produces a set of disjoint colour codes the flag graph contains unsplittable seams (corresponding to the two degree-3 vertices in the base graph). We generalise this idea with the following statement
@@
where three cycles are glued together at a set of edges. There is no ungluing that can recover these three cycles, but we can recover a pair of cycles (as shown). A flag graph obtained from three colour code lattices by a similar gluing is shown in Fig. 7 b). Even though all the seams are splittable we cannot recover these three lattices via splitting, we can only obtain a pair of lattices.

### v3

@@ II.2 The simplex graph
From Proposition˜1 we can see that an $X$ operator supported on the vertices of an $x$-rainbow subgraph commutes with all $Z$ stabilisers (and similarly for $Z$ operators on $z$-rainbow subgraphs) and so if these operators are not stabilisers they must be logicals. We will see in Section˜III that this fact is the cause of the low distances observed for the majority of pin codes studied numerically in vuillot2022quantum, as the method used to generate these codes naturally gives rise to rainbow subgraphs of size $4$. In addition, in Section˜IV we will show that it prevents useful transversal non-Clifford gates in the majority of cases.
@@ II.3 Rainbow codes
A $D$-dimensional rainbow code is a CSS code with qubits associated to the vertices of a simplex graph with $D+1$ colours and stabiliser generators associated to a subset of the $x$-maximal and $x$-rainbow subgraphs (for $X$ stabilisers) and $z$-maximal and $z$-rainbow subgraphs (for $Z$ stabilisers) of this graph, for $x$ and $z$ satisfying Eq.˜1.
@@
Note that in general mixed rainbow codes are not required to have $X$ and $Z$ stabilisers supported on all $x$- and $z$-rainbow subgraphs as such operators would not commute in general. The definitions of these different classes of rainbow code are summarised in Table˜1. Notice that all of these classes can be thought of as generalisations of colour codes as maximal and rainbow subgraphs in colour codes are equivalent. To understand how the properties of these classes can differ from each other we will need to look at some more specific examples, such as those presented in the next section. We also remark that, for all stabiliser assignments we will consider, the sets of all relevant maximal and rainbow subgraphs are efficiently computable and algorithms for finding them are presented in Appendix˜C.
@@
mixed
some $x$-maximal
and $x$-rainbow
some $z$-maximal
and $z$-rainbow
@@
Finally, we comment on the stabiliser weights of these codes. These weights are upper-bounded by the size of the largest $D$-maximal subgraph, or equivalently by the largest number of $D$-simplices sharing a common 0-simplex. In general this can be arbitrarily large but in more specific cases it can be bounded, such as in the case of the codes considered in the next section.
@@ III Rainbow codes from hypergraph products and gluings of colour codes
Given a cell complex that defines a cellulation of a manifold, a colour code on this manifold can be defined using a mapping from the chain complex to a corresponding simplicial complex bombin2007exact; zhu2023non. Similar techniques were used in vuillot2022quantum to define pin codes on more general chain complexes obtained from the hypergraph product of classical codes tillich_quantum_2014, and we could use these mappings and the results of the previous section to obtain other classes of rainbow code in an identical fashion. However, it turns out that there is an alternative perspective on the hypergraph product – and on the codes derived from it – that provides considerably more intuition into the structure and properties of these codes. In what follows we will show that any $D$-dimensional rainbow code defined via the hypergraph product can be viewed as a collection of $D$-dimensional colour codes joined together at (D1)(D-1)-dimensional domain walls. This makes understanding the code parameters and action of logical gates much more straightforward, and also implies connections between these codes and other constructions involving some forms of code gluing, such as welded codes michnicki20143d, defect networks song2023topological, and layer codes williamson_layer_2024.
We begin by studying gluing operations on graphs and their interplay with the hypergraph product in Section˜III.1. In Section˜III.2 we show how to obtain flag graphs from the output of the hypergraph product and demonstrate the effects of gluing operations on these graphs. In Section˜III.3 we show how, when all inputs to the product are cycles, the resulting flag graph can be understood as a collection of colour code lattices joined at objects we call seams, and finally we show how to assign stabilisers to these seams and study their effect on colour code logical operators in Section˜III.4.
@@ III.1 Gluings of graphs
To start, we define the hypergraph product (HGP) and show how its action commutes with gluing operations on graphs. This product is commonly defined in terms of tensor products of chain complexes but it can also be understood as a Cartesian product of bipartite graphs  [^2] If we label the two kinds of vertex type-0 and type-$1$ then the adjacency matrix with type-0 vertices as rows and type-$1$ vertices as columns is equivalent to a representation of the boundary map from $1$-cells to 0-cells in the chain complex picture. Recall that a graph $\mathcal{G}$ is a set of vertices $V_{\mathcal{G}}$ and a set of edges $E_{\mathcal{G}}$, where an edge is an (unordered) pair of vertices $\{g_{1},g_{2}\}$. Also, the Cartesian product of two graphs is defined as
@@
Given a graph $\mathcal{G}\square\mathcal{H}$, the operation $g_{1}\Leftarrow g_{2}:\mathcal{G}\square\mathcal{H}$ is a gluing $((g_{1},h_{i})\leftarrow(g_{2},h_{i}):\mathcal{G}\square\mathcal{H})\penalty 10000\ \forall\penalty 10000\ h_{i}\in V_{\mathcal{H}}$
where the order of these gluings does not matter by Lemma˜5. This operation generalises straightforwardly to gluing along a hyperplane in a graph $\mathcal{G}=\mathcal{A}\square\mathcal{B}\square$... since we can use the associativity and commutativity of the Cartesian product to write
[…]
@@
$\{\{g_{i},h_{j}\}\penalty 10000\ |\penalty 10000\ (g_{i}\neq g_{2})\in V_{\mathcal{G}},h_{j}\in V_{\mathcal{H}}\}.$ (8)
On the other hand, $g_{1}\Leftarrow g_{2}:\mathcal{G}$ results in a vertex set $V_{\mathcal{G}}^{\prime}=\{g_{i}\penalty 10000\ |\penalty 10000\ (g_{i}\neq g_{2})\in V_{\mathcal{G}}\}$ and the product $(g_{1}\Leftarrow g_{2}:\mathcal{G})\square\mathcal{H}$ then has vertex set
$\{\{g_{i},h_{j}\}\penalty 10000\ |\penalty 10000\ g_{i}\in V_{\mathcal{G}}^{\prime},h_{j}\in V_{\mathcal{H}}\}$ (9)
@@
Secondly, we can show that each vertex has the same neighbourhood in both graphs (equivalent to showing that the edge sets are equivalent). First, notice that for a vertex $(g_{i},h_{j})$ of $\mathcal{G}\square\mathcal{H}$, the neighbourhood $\mathcal{N}((g_{i},h_{j}))$ can be split into two parts, $N^{g}$ and $N^{h}$, where $N^{g}=\mathcal{N}(g_{i})\times h_{j}$ and $N^{h}=g_{i}\times\mathcal{N}(h_{j})$. Then, recalling the effect of gluing on neighbourhoods described in the proof of Lemma˜5, for $g_{1}\Leftarrow g_{2}:\mathcal{G}\square\mathcal{H}$ we have that
@@
An example of this commutation is shown in Fig.˜3.
@@
Given a vertex $g_{1}$ of a graph $\mathcal{G}$ and a set of edges $U\subseteq\mathcal{N}(g_{1})$, an ungluing $g_{1}\overset{U}{\leftrightarrow}g_{2}:\mathcal{G}$ adds a new vertex $g_{2}$ to $V_{\mathcal{G}}$ and replaces all edges $\{g_{1},g_{i}\}\in U$ with edges $\{g_{2},g_{i}\}$
@@
Given a graph $\mathcal{G}\square\mathcal{H}$ and a set $U\in\mathcal{N}(g_{1})$, the operation $g_{1}\overset{U}{\Leftrightarrow}g_{2}:\mathcal{G}\square\mathcal{H}$ is equivalent to the operation $(g_{1}\overset{U}{\leftrightarrow}g_{2}:\mathcal{G})\square\mathcal{H}$.
@@
$\begin{split}&g_{1}\Leftarrow g_{2}:(g_{1}\overset{U}{\Leftrightarrow}g_{2}:\mathcal{G}\square\mathcal{H})\\ =\penalty 10000\ &g_{1}\Leftarrow g_{2}:((g_{1}\overset{U}{\leftrightarrow}g_{2}:\mathcal{G})\square\mathcal{H})\\ =\penalty 10000\ &\mathcal{G}\square\mathcal{H}\end{split}$ (16)
is then guaranteed by Proposition˜2 and Eq.˜14.
The insight provided by Eq.˜16 that instead of studying the product of a few very complex graphs, we can equivalently “unglue” these graphs into simpler graphs, take products of these and then glue together the results. However, we still need a way to define simplicial complexes from these graphs, as well as way to understand the effect of gluing on these complexes.
@@ III.2 Flags and their gluings
More generally, for the Cartesian product of $D$ bipartite graphs we will have $D+1$ levels of vertex in the output, with the level of a given output vertex equal to the sum of the levels of all its corresponding input vertices. Typically we will be focused on a single input graph $\mathcal{G}$ (i.e. as the target of a gluing) and in this case we will use an abbreviated notation $(g_{i}^{l},...)^{L}$ where $l\in\{0,1\}$, $L$ is the level of the product vertex and the ... represents input vertices from all other input graphs. A $D$-dimensional simplicial complex can then be obtained from this graph by associating a $D$-simplex with every path of length $D+1$ that consists of a sequence of vertices with levels 0,1,,D0,1,...,D. In keeping with the terminology of vuillot2022quantum we will refer to these paths as flags and write them as $(D+1)$-uples $((g_{i}^{0},...)^{0},...,(g_{i^{\prime}}^{1},...)^{D})$. Two examples for $D=2$ and $D=3$ are shown in Fig.˜4. When discussing simplicial complexes obtained in this manner we will use the term “flag graph” (or $\mathcal{G}_{\mathcal{F}}$) to refer to the simplex graph and “product graph” (or $\mathcal{G}_{\square}$) to refer to the graph with levelled vertices obtained from the Cartesian product. Additionally, the vertex levels in the product graph will be used to label the edge colours in the flag graph, so that e.g. two flags which share product graph vertices with levels 1,2,,D1,2,...,D will correspond to flag graph vertices connected by an edge with colour $c_{0}$.
@@
Figure 5: Gluings on product graphs (numbered vertices and black edges) and their effects on the corresponding flag graphs (dark vertices and coloured edges). a) A single gluing that joins two flags at D1D-1 vertices in the product graph and so adds a new $c_{0}$-coloured edge between the corresponding vertices of the flag graph. b) A pair of gluings that glue together pairs of flags in the product graph and so glue together the corresponding vertices in the flag graph. Seams in the product graph are shown in both cases by grey edges/vertices.
[…]
@@
In the first case a new $c_{D}$ edge in the flag graph will be created between the vertices corresponding to $F_{1}$ and $F_{2}$, while in the second case a new $c_{0}$ edge will be created. No other colours of edge can be created by gluing. An example of such a gluing is shown in Fig.˜5 a). We then fix the following terminology.
@@
As alluded to above, it is possible to use multiple gluings to glue one flag fully onto another, and thus glue together a pair of vertices in the flag graph. This is accomplished using a pair of gluings $g_{i}^{0}\Leftarrow g_{j}^{0}$ and $g_{i^{\prime}}^{1}\Leftarrow g_{j^{\prime}}^{1}$ where $g_{i^{\prime}}^{1}\in\mathcal{N}(g_{i}^{0})$ and $g_{j^{\prime}}^{1}\in\mathcal{N}(g_{j}^{0})$. This creates a pair of seams, as in Fig.˜5 b).
We also care about the action of ungluing on flag graphs. From Eq.˜16 we know that $\mathcal{G}_{\square}$ and $g_{1}^{l}\Leftarrow g_{2}^{l}:(g_{1}^{l}\overset{U}{\Leftrightarrow}g_{2}^{l}:\mathcal{G}_{\square})$ are identical and so have identical flag graphs. Since $g_{1}^{l}\Leftarrow g_{2}^{l}$ adds edges of colour $c_{0}$ or $c_{D}$, $g_{1}^{l}\overset{U}{\Leftrightarrow}g_{2}^{l}$ must therefore delete edges of colour $c_{0}$ or $c_{D}$.
@@ III.3 Joining colour code lattices
Because $\mathcal{G}_{\square}$ is a cellulation of a $D$-dimensional manifold the simplicial complex described by $\mathcal{G}_{\mathcal{F}}$ is also a cellulation of this same manifold (it is simply a subdivision of the hypercubic lattice into simplices), and by Lemma˜2 $\mathcal{G}_{\mathcal{F}}$ is then a $D$-dimensional colour code lattice. ∎
@@
An example of this is shown in Fig.˜6 a). This statement is equivalent to results in other works where $\mathcal{G}_{\square}$ is interpreted as a chain complex rather than a graph bombin2007exact; vuillot2022quantum.
@@
Now consider a graph $\mathcal{G}_{\square}=(\mathcal{O}_{1}\cup\mathcal{O}_{2})\square\mathcal{O}_{3}\square...\square\mathcal{O}_{D+1}$ which is equivalent to a pair of disjoint $D$-dimensional hypercubic lattices, and so has a flag graph equivalant to a pair of disjoint $D$-dimensional colour code lattices. The gluing $o_{1,i}^{l}\Leftarrow o_{2,j}^{l}:\mathcal{G}_{\square}$ which glues together these two hypercubic lattices then also modifies the associated colour code lattices/flag graph in accordance with the discussion in the previous section. Specifically, it creates new $c_{0}$ or $c_{D}$ edges (depending on $l=1$ or 0) between vertices of $\mathcal{G}_{\mathcal{F}}$ that correspond to flags lying on the seam assocaited to the gluing, and so joins the whole flag graph into a single connected component. An example is shown in Fig.˜6 b), where we can see that the flag graph post-gluing remains locally identical to a colour code lattice everywhere except in the neighbourhood of the seam. We refer to this operation as the join of two colour code lattices.
@@
Given a pair of joined colour code lattices defined from a graph $\mathcal{G}=g_{i}^{l}\Leftarrow g_{j}^{l}:(\mathcal{G}_{\square}^{1}\cup\mathcal{G}_{\square}^{2})$, a split of these lattices is the transformation induced by the ungluing $g_{i}^{l}\overset{U}{\Leftrightarrow}g_{j}^{l}:\mathcal{G}$, where $\lvert U\rvert=2$.
Notice that, due to Eq.˜15, it is not guaranteed that performing a join and then a split recovers the original pair of colour code lattices. However, the requirement that $\lvert U\rvert=2$ ensures that we recover either a pair of colour code lattice or a single larger colour code lattice, as this ungluing necessarily divides the underlying glued cycle graphs back into a disjoint union of cycle graphs. We can also convince ourselves of this diagramatically by considering a gluing of two cycles
@@
Sets of colour code lattices that can be mapped into each other by performing a join and then a split will be referred to as split-equivalent. Examples are shown in Fig.˜6 c) and d).
@@
where two cycles are glued at three vertices/two edges. Because two of the vertices in the resulting graph have degree three there is no way to recover a pair of cycles using the ungluing operation that we have defined. We show in Fig.˜7 the effect of such a gluing on a pair of colour code lattices. Because the gluing is not reversible there is no splitting of this flag graph that produces a set of disjoint colour codes the flag graph contains unsplittable seams (corresponding to the two degree-3 vertices in the base graph). We generalise this idea with the following statement
@@
where three cycles are glued together at a set of edges. There is no ungluing that can recover these three cycles, but we can recover a pair of cycles (as shown). A flag graph obtained from three colour code lattices by a similar gluing is shown in Fig.˜7 b). Even though all the seams are splittable we cannot recover these three lattices via splitting, we can only obtain a pair of lattices.